HETAV PATEL

Grade 11,

Find the period f(x)=[sin3x]+|cos6x| Thanks

Find the period
f(x)=[sin3x]+|cos6x|

Thanks

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
f(x) = [sin(3x)] + |cos(6x)|
f(x) = f_{1}(x) + f_{2}(x)
f_{1}(x) = [sin(3x)]
= 0, 0\leq x\leq \frac{\pi }{3}
= -1, \frac{\pi }{3}\leq x\leq \frac{2\pi }{3}
Time period of f1(x):
\frac{2\pi }{3} = \frac{4\pi }{6}
f_{2}(x) = |cos(6x)|
Time period of f2(x):
\frac{\pi }{6}
Time period of f(x):
(LCM of time period of f1(x) & f2(x))/(HCFof time period of f1(x) & f2(x))
\frac{4\pi }{6} = \frac{2\pi }{3}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty